Consider the following production function: Q=(K1/2+L1/2)2.

  • Find the cost (wL+rK) minimizing quantities of capital (K) and labor (L) to produce a given quantity Q, i.e. minimize wL+rK s.t. Q=(K1/2+L1/2)2.
  • Check SOCs for a minimum.
Answer #1

Q) Q = (kV2 + 118 )?, at Cost Minimizations MRTS LK = Wlr . MRTS LK = MPL = 2 ( SK + FL) (Vg L-1/2) = 1 kl. MPAS 2 (se) (Vgraye No. NOW SOC 9- I as .L = WLtok taco- (SK +$2) ?] poc. aL - W - X 2 (5k +52) 1 =0. I OL odpis = 7-ala (Tat -0. 80.500 h =